OpenStudy (anonymous):

in(x-6)=1 SHOW work !PLEASE !!!!

OpenStudy (anonymous):

\[\ln(x-6)=1\] rewrite in exponential form as \[e^{1}=x-6\] or \[e=x-6\] and solve for \(x\)

OpenStudy (anonymous):

\[\log_{10}(m) = n \Rightarrow m = (10)^n \]

OpenStudy (anonymous):

you should be able to go from \[b^x=y\] to \[\log_b(y)=x\] and back easily

OpenStudy (anonymous):

\[(x-6) = (10)^1 \Rightarrow x = 10 + 6 = 16 \]

OpenStudy (anonymous):

btw if you see \(\ln(x)\) you visualize it at \(\log_e(x)\)

OpenStudy (anonymous):

So my answer would be : 16 ?

terenzreignz (terenzreignz):

No... follow up from sat... e = x - 6 and solve for x. Add 6 to both sides, and that should do it.

OpenStudy (anonymous):

ln(x-6)=1 or we can write like this lnx/ln6=1 =>lnx=ln6 =>x=6